A company is projected to have a free cash flow of $336 million next year, growing at a 6% rate until the end of year 3. After that, cash flows are expected to grow at a stable rate of 2.3% in perpetuity. The company's cost of capital is 9.7%. The company owes $89 million to lenders and has $18 million in cash. If it has 186 million shares outstanding, what is your estimate for its stock price? Round to one decimal place. (e.g., $4.32 = 4.3)
Answer #1

Firm Value = [FCF1 / (1 + r)] + [{FCF1 * (1 + g1)} / (1 + r)2] + [{FCF1 * (1 + g1)2} / (1 + r)3] + [{FCF1 * (1 + g1)2 * (1 + gC)} / {(r - gC) * (1 + r)3}]

= [336 / (1 + 0.097)] + [{336 * (1 + 0.06)} / (1 + 0.097)2] + [{336 * (1 + 0.06)2} / (1 + 0.097)3] + [{336 * (1 + 0.06)2 * (1 + 0.023)} / {(0.097 - 0.023) * (1 + 0.097)3}]

= 306.29 + 295.96 + 285.98 + 3,953.44 = $4,841.67 million

Equity Value = Firm Value - Debt Value + Cash

= 4,841.67 - 89 + 18 = $4,770.67 million

Share Price = Equity Value / No. of shares outstanding = 4,770.67 / 186 = $25.6